icu lact that thie muiiplication of matrices with real coefficients is associative. 1. Show that every cyclic group is abelian. 2. Let m be an integer greater than 1. For any l e Z, let Im denote the remain- der of the division of I by m. Let: Zm = {0. 1, 2, . . . ,m-1)- on Zm as follows: Define a binary operation where + is the usual addition of integers. (a) Show that G (Zm,*) is a group. (b) Show that for any j E G, we have: TT7L ord j ged(m,J) (c) Show that G = (225, *) is cyclic, and find all of its generators. 3·Let SL(2, Z) denote the set of 2 x 2 matrices with integer coefficients whose matrix multiplica- determinants are equal to 1. Is SL(2,Z) a group under



I am not sure how to do
Answer #1

Answer-1. Let G be a cyclic group. Then, there exists a\in G such that G=<a>=\{a^m:m\in\mathbb{Z}\}. Now, let x,y\in G. Then, there exist m,n\in \mathbb{Z} such that x=a^m and y=a^n. Now,xy=a^ma^n=a^{m+n}=a^{n+m}=a^na^m=yx and hence it follows that for any x,y\in G,xy=yx, that is, the cyclic group G is abelian.